OpenStudy (anonymous):

Find the quotient of 3/7 & 5/8

jhonyy9 (jhonyy9):

3/7 --- 0,4 5/8 --- 0,6

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\frac{3}{7}\div \frac{5}{8}=\frac{3}{7}\times \frac{8}{5}\]